OpenStudy (anonymous):

What is the equation of a circle with the center of -2,1 and passes through 3,4

OpenStudy (loser66):

what is the formula of a circle?

OpenStudy (jdoe0001):

 you have the center it passes through 3,4 thus find the radius then recall that \(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{purple}{ r}}\)

OpenStudy (jdoe0001):

and thus \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -2}}\quad ,&{\color{blue}{ 1}})\quad % (c,d) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

So would it be (x-2)^2 + (x+1) =13

OpenStudy (anonymous):



OpenStudy (anonymous):

then the equation of circle is (x+2)^2+(y+1)^2=r^2=50 plus because centre is (-2,-1) remember( -*-)=+